# Hurwitz moduli varieties parameterizing pointed covers of an algebraic curve with a fixed monodromy group

### v3

@@ 4. Universal families of pointed covers with fixed monodromy group
For every $\sigma\in N(\lambda_{0})$ one has $\sigma G(\lambda_{0})\sigma^{-1}=G(\lambda_{0})$, so $\Gamma_{\sigma m\sigma^{-1},\lambda_{0}}=\Gamma_{m,\lambda_{0}}$. A basis of open sets of the topology of ${\mathcal{C}(y_{0})^{\prime}}^{an}$ was constructed in [21, Section 3] as follows. Let $(y,(D,m))\in Y\times H^{G}_{n}(Y,y_{0})\setminus B$, $D=\{b_{1},\ldots,b_{n}\}$. Let $\overline{U},\overline{U}_{1},\ldots,\overline{U}_{n}$ be disjoint embedded closed disks in $Y$ with interiors $U,U_{1},\ldots,U_{n}$ respectively, such that $y\in U$, $\overline{U}_{i}\subset Y\setminus\{y_{0}\}$, $b_{i}\in U_{i}$, $i=1,\ldots,n$. Let $\alpha:I\to Y\setminus\cup_{i=1}^{n}\overline{U}_{i}$ be a path such that $\alpha(0)=y_{0}$, $\alpha(1)=y$. Let
@@
(21) ${p^{\prime}}^{-1}(U\times N_{(D,m)}(U_{1},\ldots,U_{n}))=\bigsqcup_{j=1}^{|G|}N_{(\alpha_{j},D,m)}(U,U_{1},\ldots,U_{n})$
@@
(22) $\begin{split}\overline{N}_{(\alpha,D,m)}(U,U_{1},\ldots,U_{n})=\{&(\Gamma_{m(E),\lambda_{0}}[\alpha\cdot\tau]_{E},E,m(E))|\\ &E\in N_{D}(U_{1},\ldots,U_{n}),\tau:I\to U,\tau(0)=y\}\end{split}$
@@
${\varphi^{\prime}}^{-1}(U\times N_{(D,m)}(U_{1},\ldots,U_{n}))=\bigsqcup_{i=1}^{d}\overline{N}_{(\alpha_{i},D,m)}(U,U_{1},\ldots,U_{n})$
@@
By [21, Corollary 4.5] $N(\lambda_{0})$ acts on $\mathcal{X}(y_{0})^{\prime}$ by covering automorphisms of the composed étale cover (19). It is clear from (18) that $\varphi^{\prime}$ is $N(\lambda_{0})$-equivariant.
Let $H$ be a connected component of $H^{G}_{n}(Y,y_{0})$. Let $\mathcal{X}(y_{0})^{\prime}_{H}={\varphi^{\prime}}^{-1}(Y\times H\setminus B)$. This is an irreducible algebraic variety, quotient of the irreducible variety ${p^{\prime}}^{-1}(Y\times H\setminus B)=\mathcal{C}(y_{0})^{\prime}_{H}$ (cf. [21, § 3.16]) and the smooth variety $\mathcal{X}(y_{0})_{H}=\linebreak\varphi^{-1}(Y\times H)$ is the normalization of $Y\times H$ in the field of rational functions $\mathbb{C}(\mathcal{X}(y_{0})^{\prime}_{H})$. Given $\sigma\in N(\lambda_{0})$, the map $x\mapsto\sigma x$ defines an isomorphism $\mathcal{X}(y_{0})^{\prime}_{H}\overset{\sim}{\longrightarrow}\mathcal{X}(y_{0})^{\prime}_{\sigma\ast H}$. Passing to normalizations this isomorphism extends in a unique way to an isomorphism $\mathcal{X}(y_{0})_{H}\overset{\sim}{\longrightarrow}\mathcal{X}(y_{0})_{\sigma\ast H}$. This defines an action of $N(\lambda_{0})$ on $\mathcal{X}(y_{0})$ with the required properties. ∎
@@
The group $N(\lambda_{0})$ acts on $\mathcal{X}(y_{0})$ by covering automorphisms of the cover $\mathcal{X}(y_{0})\to Y\times(Y\setminus y_{0})^{(n)}_{\ast}$. We denote by $\mathcal{X}(y_{0},\lambda_{0})$ the quotient algebraic variety $\mathcal{X}(y_{0})/N(\lambda_{0})$. The $N(\lambda_{0})$-equivariant cover $\varphi:\mathcal{X}(y_{0})\to Y\times H^{G}_{n}(Y,y_{0})$ descends to a cover $\phi:\mathcal{X}(y_{0},\lambda_{0})\to Y\times H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$. One has the following commutative diagram
(23) ${\mathcal{X}(y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\kappa}$$\scriptstyle{\varphi}$${\mathcal{X}(y_{0},\lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${Y\times H^{G}_{n}(Y,y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{id\times\nu}$${Y\times H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})}$
where $\kappa$ and $\nu$ are the quotient morphisms with respect to the action of $N(\lambda_{0})$. The morphism $\kappa\circ\eta:H^{G}_{n}(Y,y_{0})\to\mathcal{X}(y_{0},\lambda_{0})$ maps $(D,m)$ to $(\Gamma_{m,\lambda_{0}}[c_{y_{0}}]_{D},D,m^{N(\lambda_{0})})$, where $m^{N(\lambda_{0})}=\{\sigma m\sigma^{-1}|\sigma\in N(\lambda_{0})\}$, so $\kappa\circ\eta$ is $N(\lambda_{0})$-invariant and can be decomposed as $\xi\circ\nu$, where $\xi:H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})\to\mathcal{X}(y_{0},\lambda_{0})$ is a morphism with the property that $\phi(\xi(D,m^{N(\lambda_{0})}))=(y_{0},(D,m^{N(\lambda_{0})}))$ for $\forall(D,m^{N(\lambda_{0})})\in H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$.
The closed algebraic subset $B\subset Y\times H^{G}_{n}(Y,y_{0})$ is $N(\lambda_{0})$-invariant. Let $\mathcal{B}=\linebreak id\times\nu(B)$ and let $\mathcal{X}(y_{0},\lambda_{0})^{\prime}=\mathcal{X}(y_{0},\lambda_{0})\setminus\phi^{-1}(\mathcal{B})$. Then $\mathcal{X}(y_{0},\lambda_{0})^{\prime}=\mathcal{X}(y_{0})^{\prime}/N(\lambda_{0})$ and $|{\mathcal{X}(y_{0},\lambda_{0})^{\prime}}^{an}|=|{\mathcal{X}(y_{0})^{\prime}}^{an}|/N(\lambda_{0})$ (cf. [21, Lemma 2.5]). For every $(y,(D,m^{N(\lambda_{0})}))\in Y\times H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})\setminus\mathcal{B}$ and every $z=(\Gamma_{m,\lambda_{0}}[\alpha]_{D},D,m^{N(\lambda_{0})})\in\mathcal{X}(y_{0},\lambda_{0})^{\prime}$ such that $\phi(z)=(y,(D,m^{N(\lambda_{0})}))$, the sets
@@
The map $\beta:S\to(Y\setminus y_{0})^{(n)}_{\ast}\subset Y^{(n)}$ given by $\beta(s)=B_{s}$ is a morphism by [21, Proposition 2.6]. The map $u$ fits in the following commutative diagram
@@
where $\delta_{1}$ is the étale cover defined by $\delta_{1}(D,m^{N(\lambda_{0})})=D$ (cf. Proposition 4.8). In order to prove that $u$ is a morphism it suffices to verify that $u$ is continuous with respect to the topologies of $S^{an}$ and $H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})^{an}$ (cf. [21, Corollary 4.5]), i.e. we have to show that for every $s\in S$ and every neighborhood $N$ of $u(s)$ the point $s$ is internal of $u^{-1}(N)$. Let $s_{0}$ be an arbitrary point of $S$. Let $\beta(s_{0})=D=\{b_{1},\ldots,b_{n}\}$. Let ($\varepsilon:\Lambda\to f_{s_{0}}^{-1}(y_{0})$, $m:\pi_{1}(Y\setminus D,y_{0})\to G$) be a pair which satisfies Condition (i) of Proposition 2.8 with $\varepsilon(\lambda_{0})=\eta(s_{0})$. One has $u(s_{0})=(D,m^{N(\lambda_{0})})$. Let $N_{(D,m^{N(\lambda_{0})})}(U_{1},\ldots,U_{n})$ be any of the open sets of the neighborhood basis of $(D,m^{N(\lambda_{0})})$ in $|H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})^{an}|$ contained in $N$ (cf. (16)). In the course of the proof of Lemma 3.5 we showed that there exists a connected neighborhood $V\subset|S^{an}|$ of $s_{0}$ such that $\beta(V)\subset N_{D}(U_{1},\ldots,U_{n})$ and we constructed for every $s\in V$ a pair $(\varepsilon^{\prime}_{s}:\Lambda\to f^{-1}_{s}(y_{0}),m^{\prime}_{s}=m(\beta(s)))$ as in (10). For $\forall s\in S$ this pair satisfies the conditions of Proposition 2.8(i). In fact, the first one was verified in the proof of Lemma 3.5. We claim that the second one is satisfied as well: $\varepsilon^{\prime}_{s}(\lambda_{0})=\eta(s)$ for $\forall s\in S$. This holds since $\{y_{0}\}\times V$ is a connected subset of $U\times V$, so $f^{-1}(\{y_{0}\}\times V)$ is a disjoint union of $d=|\Lambda|$ connected components $f^{-1}(\{y_{0}\}\times V)\cap W_{\lambda}$, $\lambda\in\Lambda$. The map $\eta:S\to f^{-1}(\{y_{0}\}\times S)$ is continuous with respect to the canonical complex topologies and $\eta(s_{0})=\varepsilon(\lambda_{0})\in W_{\lambda_{0}}$. Therefore $\eta(V)$ is the connected component of $f^{-1}(\{y_{0}\times V\})$ contained in $W_{\lambda_{0}}$. This shows that $\varepsilon_{s}(\lambda_{0})=\eta(s)$ for $\forall s\in V$. Using Proposition 2.8 we conclude that for $\forall s\in V$ the monodromy invariant $u(s)$ of $(f_{s}:X_{s}\to Y,\eta(s))$ equals $(\beta(s),m(\beta(s))^{N(\lambda_{0})})$, so $u(s)\in\nu(N_{(D,m)}(U_{1},\ldots,U_{n}))=N_{(D,m^{N(\lambda_{0})})}(U_{1},\ldots,U_{n})$ (cf. (16)). Therefore $V\subset u^{-1}(N)$. ∎
@@
Let $f:X\to S$ and $g:Z\to S$ be proper morphisms of algebraic varieties. Let $h:X\to Z$ be a morphism such that $f=g\circ h$. Suppose that for every $s\in f(X)$ the induced morphism of the scheme-theoretical fibers $h_{s}:X\otimes_{S}\mathbb{C}(s)\to Z\otimes_{S}\mathbb{C}(s)$ is an isomorphism. Then $h$ is an isomorphism.
@@
The map $h$ is bijective. By [14, Proposition 4.6.7(i)] every $s\in f(X)$ has an open neighborhood $U$ in $S$ such that $h|_{f^{-1}(U)}:f^{-1}(U)\to g^{-1}(U)$ is a closed embedding. Since $X$ and $Z$ are reduced schemes $h|_{f^{-1}(U)}$ is an isomorphism. Therefore $h$ is an isomorphism. ∎
In the next theorem we assume the set-up of § 2.1.
@@
The algebraic variety $H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$ is a fine moduli variety for the moduli functor ($\mathcal{H}^{\Lambda,G}_{(Y,y_{0}),n}$ of smooth, proper families of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ branched in $n$ points (cf. § 3.8). The universal family is (cf. Theorem 4.13)
@@
Let $[f:X\to Y\times S,\eta:S\to X]\in\mathcal{H}^{\Lambda,G}_{(Y,y_{0}),n}(S)$ and let $B\subset Y\times S$ be the branch locus of $f$. For every fiber ($f_{s}:X_{s}\to Y,\eta(s))$ let $(\varepsilon_{s},m_{s})$ be a pair which satisfies the conditions of Proposition 2.8(i) (cf. Proposition 3.3). Let $u:S\to H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$, $u(s)=(B_{s},m_{s}^{N(\lambda_{0})})$ be the morphism of Proposition 4.15. We want to prove that $(f:X\to Y\times S,\eta)$ is equivalent to the pullback by $u$ of the family (26). This is the unique morphism with this property since the monodromy invariant classifies the pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ up to equivalence. For every $s\in S$ there exists a unique isomorphism $h_{s}:X_{s}\to\mathcal{X}(y_{0},\lambda_{0})_{u(s)}$ such that $\phi_{u(s)}\circ h_{s}=(id_{Y}\times u)\circ f_{s}$ and $h_{s}(\eta(s))=\xi(u(s))$. Let $h:X\to\mathcal{X}(y_{0},\lambda_{0})$ be the map whose restriction on every $X_{s}$ equals $h_{s}$. One obtains the following commutative diagram of maps
(27) ${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$$\scriptstyle{f}$${\mathcal{X}(y_{0},\lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${Y\times S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{id\times u}$${Y\times H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})}$
@@
By the covering homotopy property of topological covering maps (cf. [29, Ch.2 § 3 Th.3]) there is a unique continuous lifting of $F$
[…]
@@
One applies [21, Corollary 4.5] to the commutative diagram
@@
Let $\phi_{S}:\mathcal{X}(y_{0},\lambda_{0})_{S}\to Y\times S$ be the pullback of $\phi:\mathcal{X}(y_{0},\lambda_{0})\to Y\times H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$ by $u:S\to H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$. The composition $X^{\prime}\overset{h^{\prime}}{\longrightarrow}\mathcal{X}(y_{0},\lambda_{0})^{\prime}\hookrightarrow\mathcal{X}(y_{0},\lambda_{0})$ yields an $S$-morphism $\psi:X^{\prime}\to\mathcal{X}(y_{0},\lambda_{0})_{S}$ which fits in the following commutative diagram of morphisms

### v4

@@ 4. Universal families of pointed covers with fixed monodromy group
For every $\sigma\in N(\lambda_{0})$ one has $\sigma G(\lambda_{0})\sigma^{-1}=G(\lambda_{0})$, so $\Gamma_{\sigma m\sigma^{-1},\lambda_{0}}=\Gamma_{m,\lambda_{0}}$. A basis of open sets of the topology of ${\mathcal{C}(y_{0})^{\prime}}^{an}$ was constructed in [22, Section 3] as follows. Let $(y,(D,m))\in Y\times H^{G}_{n}(Y,y_{0})\setminus B$, $D=\{b_{1},\ldots,b_{n}\}$. Let $\overline{U},\overline{U}_{1},\ldots,\overline{U}_{n}$ be disjoint embedded closed disks in $Y$ with interiors $U,U_{1},\ldots,U_{n}$ respectively, such that $y\in U$, $\overline{U}_{i}\subset Y\setminus\{y_{0}\}$, $b_{i}\in U_{i}$, $i=1,\ldots,n$. Let $\alpha:I\to Y\setminus\cup_{i=1}^{n}\overline{U}_{i}$ be a path such that $\alpha(0)=y_{0}$, $\alpha(1)=y$. Let
@@
(21) ${p^{\prime}}^{-1}(U\times N_{(D,m)}(U_{1},\ldots,U_{n}))=\bigsqcup_{j=1}^{|G|}N_{(\alpha_{j},D,m)}(U,U_{1},\ldots,U_{n}),$
@@
(22) $\begin{split}\overline{N}_{(\alpha,D,m)}(U,U_{1},\ldots,U_{n})=\{&(\Gamma_{m(E),\lambda_{0}}[\alpha\cdot\tau]_{E},E,m(E))|\\ &E\in N_{D}(U_{1},\ldots,U_{n}),\tau:I\to U,\tau(0)=y\}.\end{split}$
@@
${\varphi^{\prime}}^{-1}(U\times N_{(D,m)}(U_{1},\ldots,U_{n}))=\bigsqcup_{i=1}^{d}\overline{N}_{(\alpha_{i},D,m)}(U,U_{1},\ldots,U_{n}),$
@@
By [22, Corollary 4.5] $N(\lambda_{0})$ acts on $\mathcal{X}(y_{0})^{\prime}$ by covering automorphisms of the composed étale cover (19). It is clear from (18) that $\varphi^{\prime}$ is $N(\lambda_{0})$-equivariant.
Let $H$ be a connected component of $H^{G}_{n}(Y,y_{0})$. Let $\mathcal{X}(y_{0})^{\prime}_{H}={\varphi^{\prime}}^{-1}(Y\times H\setminus B)$. This is an irreducible algebraic variety, quotient of the irreducible variety ${p^{\prime}}^{-1}(Y\times H\setminus B)=\mathcal{C}(y_{0})^{\prime}_{H}$ (cf. [22, §3.16]) and the smooth variety $\mathcal{X}(y_{0})_{H}=\linebreak\varphi^{-1}(Y\times H)$ is the normalization of $Y\times H$ in the field of rational functions $\mathbb{C}(\mathcal{X}(y_{0})^{\prime}_{H})$. Given $\sigma\in N(\lambda_{0})$, the map $x\mapsto\sigma x$ defines an isomorphism $\mathcal{X}(y_{0})^{\prime}_{H}\overset{\sim}{\longrightarrow}\mathcal{X}(y_{0})^{\prime}_{\sigma\ast H}$. Passing to normalizations this isomorphism extends in a unique way to an isomorphism $\mathcal{X}(y_{0})_{H}\overset{\sim}{\longrightarrow}\mathcal{X}(y_{0})_{\sigma\ast H}$. This defines an action of $N(\lambda_{0})$ on $\mathcal{X}(y_{0})$ with the required properties. ∎
@@
The group $N(\lambda_{0})$ acts on $\mathcal{X}(y_{0})$ by covering automorphisms of the cover $\mathcal{X}(y_{0})\to Y\times(Y\setminus y_{0})^{(n)}_{\ast}$. We denote by $\mathcal{X}(y_{0},\lambda_{0})$ the quotient algebraic variety $\mathcal{X}(y_{0})/N(\lambda_{0})$. The $N(\lambda_{0})$-equivariant cover $\varphi:\mathcal{X}(y_{0})\to Y\times H^{G}_{n}(Y,y_{0})$ descends to a cover $\phi:\mathcal{X}(y_{0},\lambda_{0})\to Y\times H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$. One has the following commutative diagram:
(23) ${\mathcal{X}(y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\kappa}$$\scriptstyle{\varphi}$${\mathcal{X}(y_{0},\lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${Y\times H^{G}_{n}(Y,y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{id\times\nu}$${Y\times H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0}),}$
where $\kappa$ and $\nu$ are the quotient morphisms with respect to the actions of $N(\lambda_{0})$. The morphism $\kappa\circ\eta:H^{G}_{n}(Y,y_{0})\to\mathcal{X}(y_{0},\lambda_{0})$ maps $(D,m)$ to $(\Gamma_{m,\lambda_{0}}[c_{y_{0}}]_{D},D,m^{N(\lambda_{0})})$, where $m^{N(\lambda_{0})}=\{\sigma m\sigma^{-1}|\sigma\in N(\lambda_{0})\}$, so $\kappa\circ\eta$ is $N(\lambda_{0})$-invariant and can be decomposed as $\xi\circ\nu$, where $\xi:H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})\to\mathcal{X}(y_{0},\lambda_{0})$ is a morphism with the property that $\phi(\xi(D,m^{N(\lambda_{0})}))=(y_{0},(D,m^{N(\lambda_{0})}))$ for $\forall(D,m^{N(\lambda_{0})})\in H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$.
The closed algebraic subset $B\subset Y\times H^{G}_{n}(Y,y_{0})$ is $N(\lambda_{0})$-invariant. Let $\mathcal{B}=\linebreak id\times\nu(B)$ and let $\mathcal{X}(y_{0},\lambda_{0})^{\prime}=\mathcal{X}(y_{0},\lambda_{0})\setminus\phi^{-1}(\mathcal{B})$. Then $\mathcal{X}(y_{0},\lambda_{0})^{\prime}=\mathcal{X}(y_{0})^{\prime}/N(\lambda_{0})$ and $|{\mathcal{X}(y_{0},\lambda_{0})^{\prime}}^{an}|=|{\mathcal{X}(y_{0})^{\prime}}^{an}|/N(\lambda_{0})$ (cf. [22, Lemma 2.5]). For every $(y,(D,m^{N(\lambda_{0})}))\in Y\times H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})\setminus\mathcal{B}$ and every $z=(\Gamma_{m,\lambda_{0}}[\alpha]_{D},D,m^{N(\lambda_{0})})\in\mathcal{X}(y_{0},\lambda_{0})^{\prime}$ such that $\phi(z)=(y,(D,m^{N(\lambda_{0})}))$, the sets
@@
The map $\beta:S\to(Y\setminus y_{0})^{(n)}_{\ast}\subset Y^{(n)}$ given by $\beta(s)=B_{s}$ is a morphism by [22, Proposition 2.6]. The map $u$ fits in the following commutative diagram:
@@
where $\delta_{1}$ is the étale cover defined by $\delta_{1}(D,m^{N(\lambda_{0})})=D$ (cf. Proposition 4.8). In order to prove that $u$ is a morphism it suffices to verify that $u$ is continuous with respect to the topologies of $S^{an}$ and $H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})^{an}$ (cf. [22, Corollary 4.5]), i.e. we have to show that for every $s\in S$ and every neighborhood $N$ of $u(s)$ the point $s$ is internal of $u^{-1}(N)$. Let $s_{0}$ be an arbitrary point of $S$. Let $\beta(s_{0})=D=\{b_{1},\ldots,b_{n}\}$. Let ($\varepsilon:\Lambda\to f_{s_{0}}^{-1}(y_{0})$, $m:\pi_{1}(Y\setminus D,y_{0})\to G$) be a pair which satisfies Condition (i) of Proposition 2.8 with $\varepsilon(\lambda_{0})=\eta(s_{0})$. One has $u(s_{0})=(D,m^{N(\lambda_{0})})$. Let $N_{(D,m^{N(\lambda_{0})})}(U_{1},\ldots,U_{n})$ be any of the open sets of the neighborhood basis of $(D,m^{N(\lambda_{0})})$ in $|H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})^{an}|$ contained in $N$ (cf. (16)). In the course of the proof of Lemma 3.5 we showed that there exists a connected neighborhood $V\subset|S^{an}|$ of $s_{0}$ such that $\beta(V)\subset N_{D}(U_{1},\ldots,U_{n})$ and we constructed for every $s\in V$ a pair $(\varepsilon^{\prime}_{s}:\Lambda\to f^{-1}_{s}(y_{0}),m^{\prime}_{s}=m(\beta(s)))$ as in (10). For $\forall s\in S$ this pair satisfies the conditions of Proposition 2.8(i). In fact, the first one was verified in the proof of Lemma 3.5. We claim that the second one is satisfied as well: $\varepsilon^{\prime}_{s}(\lambda_{0})=\eta(s)$ for $\forall s\in S$. This holds since $\{y_{0}\}\times V$ is a connected subset of $U\times V$, so $f^{-1}(\{y_{0}\}\times V)$ is a disjoint union of $d=|\Lambda|$ connected components $f^{-1}(\{y_{0}\}\times V)\cap W_{\lambda}$, $\lambda\in\Lambda$. The map $\eta:S\to f^{-1}(\{y_{0}\}\times S)$ is continuous with respect to the canonical complex topologies and $\eta(s_{0})=\varepsilon(\lambda_{0})\in W_{\lambda_{0}}$. Therefore $\eta(V)$ is the connected component of $f^{-1}(\{y_{0}\times V\})$ contained in $W_{\lambda_{0}}$. This shows that $\varepsilon_{s}(\lambda_{0})=\eta(s)$ for $\forall s\in V$. Using Proposition 2.8 we conclude that for $\forall s\in V$ the monodromy invariant $u(s)$ of $(f_{s}:X_{s}\to Y,\eta(s))$ equals $(\beta(s),m(\beta(s))^{N(\lambda_{0})})$, so $u(s)\in\nu(N_{(D,m)}(U_{1},\ldots,U_{n}))=N_{(D,m^{N(\lambda_{0})})}(U_{1},\ldots,U_{n})$ (cf. (16)). Therefore $V\subset u^{-1}(N)$. ∎
@@
Let $f:X\to S$ and $g:Z\to S$ be proper, surjective morphisms of algebraic varieties. Let $h:X\to Z$ be a morphism such that $f=g\circ h$. Suppose that for every $s\in S$ the induced morphism of the scheme-theoretical fibers $h_{s}:X\otimes_{S}\mathbb{C}(s)\to Z\otimes_{S}\mathbb{C}(s)$ is an isomorphism. Then $h$ is an isomorphism.
@@
The map $h$ is bijective. By [15, Proposition 4.6.7(i)] every $s\in S$ has an open neighborhood $U$ in $S$ such that $h|_{f^{-1}(U)}:f^{-1}(U)\to g^{-1}(U)$ is a closed embedding. Since $X$ and $Z$ are reduced schemes $h|_{f^{-1}(U)}$ is an isomorphism. Therefore $h$ is an isomorphism. ∎
In the next theorem we assume the setup of §2.1.
@@
The algebraic variety $H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$ is a fine moduli variety for the moduli functor $\mathcal{H}^{\Lambda,G}_{(Y,y_{0}),n}$ of smooth, proper families of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ branched in $n$ points (cf. §3.8). The universal family is (cf. Theorem 4.13)
@@
Let $f:X\to Y\times S,\eta:S\to X]\in\mathcal{H}^{\Lambda,G}_{(Y,y_{0}),n}(S)$ and let $B\subset Y\times S$ be the branch locus of $f$. For every fiber ($f_{s}:X_{s}\to Y,\eta(s))$ let $(\varepsilon_{s},m_{s})$ be a pair which satisfies the conditions of Proposition 2.8(i) (cf. Proposition 3.3). Let $u:S\to H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$, $u(s)=(B_{s},m_{s}^{N(\lambda_{0})})$ be the morphism of Proposition 4.15. We want to prove that $(f:X\to Y\times S,\eta)$ is equivalent to the pullback by $u$ of the family (26). This is the unique morphism with this property since the monodromy invariant classifies the pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ up to equivalence. For every $s\in S$ there exists a unique isomorphism $h_{s}:X_{s}\to\mathcal{X}(y_{0},\lambda_{0})_{u(s)}$ such that $\phi_{u(s)}\circ h_{s}=(id_{Y}\times u)\circ f_{s}$ and $h_{s}(\eta(s))=\xi(u(s))$. Let $h:X\to\mathcal{X}(y_{0},\lambda_{0})$ be the map whose restriction on every $X_{s}$ equals $h_{s}$. One obtains the following commutative diagram of maps:
[…]
@@
By the covering homotopy property of topological covering maps (cf. [30, Ch.2 §3 Th.3]) there is a unique continuous lifting of $F$
${X^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$${[0,1]\times V\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{F}}$$\scriptstyle{F}$${Y\times S\setminus B}$
@@
One applies [22, Corollary 4.5] to the commutative diagram
@@
Let $\phi_{S}:\mathcal{X}(y_{0},\lambda_{0})_{S}\to Y\times S$ be the pullback of $\phi:\mathcal{X}(y_{0},\lambda_{0})\to Y\times H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$ by $u:S\to H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$. The composition $X^{\prime}\overset{h^{\prime}}{\longrightarrow}\mathcal{X}(y_{0},\lambda_{0})^{\prime}\hookrightarrow\mathcal{X}(y_{0},\lambda_{0})$ yields an $S$-morphism $\psi:X^{\prime}\to\mathcal{X}(y_{0},\lambda_{0})_{S}$ which fits in the following commutative diagram of morphisms: